This Daily is being re-sent since it was "returned as undeliverable" this morning. ________________________________________________________ From: Wolf_Charlie on Mon, Aug 28, 1995 10:39 AM Subject: #1463 DAILY REPORT 8/28/95 To: a1; a2; b1; burch; c1; c2; c5; c6; c7; d2; d4; f1; h2; h8; h9; hstbb1; k2; l2; l3; l4; l5; Leibee_Jack; m4; m5; m7; n1; o1; p4; p5; r1; r4; r5; r6; r7; s1; s5; t1; w1; w3; w5; w6 HUBBLE SPACE TELESCOPE DAILY REPORT #1463 PERIOD COVERED: 0000Z (UTC) 08/25/95 - 0000Z (UTC) 08/28/95 Daily Status Report as of 240/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5957 (Snapshot of 3CR Radio Galaxies) The WF/PC-2 was used to make observations for radio galaxy 3C17.0 to study the morphology of the radio galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 6128 (Spatial Structure of the High Excitation Herbig-Haro Object HH80/81) The WF/PC-2 was used to make narrow band observations of the high-excitation Herbig-Haro Object HH81 to study its spatial structure. The objective is to directly observe the stratification of the high and low excitation lines on the HH object. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of FOS/BL 6166 (FOS Cycle 5: Location of Spectra and Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make wavelength calibration images to measure the stability of the FOS Blue Side wavelength scale and spectral locations. A Taled calibration observation was also taken. The observations were completed as planned, and no anomalies were reported. 1.4 Completed Two Sets of WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration images. Both photometric and narrow-band filters were used. The observations were executed as planned, and no problems were reported. 1.5 Completed FOS/RD 6093 (Very High Density Clumps and Outflowing Winds in QSO Emission Line Regions) The Faint Object Spectrograph (Red Side) was used to make ultraviolet spectrographic observations of emission line regions of quasi-stellar object Q0207-398 to study its very high density clumps and outflowing winds. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 5971 (HST Medium Deep Survey) The WF/PC-2 was used to make a long exposure medium deep sky survey observations at latitudes larger than 10 degrees taken in parallel with the prime science observations. The observations were executed as scheduled, and no anomalies were noted 1.7 Completed WF/PC-2 5930 (WF/PC-2 Imaging of Intermediate Redshift Radio Galaxies.) The WF/PC-2 was used to make observations of the intermediate redshift radio galaxy 3C330.0. The target is known to exhibit significant structure on the sub-arcsecond scale. Both line and continuum filters were used. The observations were completed as planned, and no anomalies were reported. 1.8 Completed Three Sets of WF/PC-2 6183 (WF/PC-2 Cycle 5 Calibration: Decon) The WF/PC-2 was used to make some internal flat, K-Spot, and dark calibration images. The purpose is to define decontamination procedures to be run during Cycle 5. The calibration observations were completed as scheduled, and no anomalies were reported. 1.9 Completed WF/PC-2 6133 (The Most Massive Stars in the Magellanic Clouds) The WF/PC-2 was used to make observations of star SK-70D69 which is known to be one of the most massive stars in the Magellanic clouds (maybe up to 200 solar masses). It is not clear whether or not these objects are single stars. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.10 Completed GHRS/1 5417 (A Search for the Physical Processes on Jupiter) The High Resolution Spectrograph (Side One) was used as part of a program to observe atomic hydrogen emission at different latitude and longitudes on the planet Jupiter. Wave calibration and sky background observations were made. The observations were completed as planned, and no anomalies were reported. 1.11 Completed FOS/RD 5607 (Probing Excitation Mechanisms and Abundances in Young Magellanic Cloud) The Faint Object Spectrograph (Red Side) was used to make spectrographic observations of supernova remnant N132D-P3 which is located in the Large Magellanic Cloud. This is to obtain spectra of three individual knots in the remnant to test emission mechanisms and derive elemental abundances. The observations were completed as planned, and no problems were reported. 1.12 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking ultraviolet observations running concurrently with prime science observations. The intent is to search for extra-galactic sources with flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The observations were completed as planned, and no problems were reported. 1.13 Completed WF/PC-2 6119 (Snapshot Survey for Companions of Planetary Nebula Nuclei) The WF/PC-2 was used to make snapshot survey observations of the planetary nebula NGC-6309 to study its nucleus and to look for possible companions. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 5972 (Determination of the Extragalactic Distance Scale) The WF/PC-2 was used to make observations of the galaxy NGC-1365 looking for standard candle Cepheid variable stars. The purpose of this observations is to help estimate the Hubble constant with an accuracy of 10%. The observations were executed as scheduled, and no anomalies were noted. 1.15 Competed WF/PC-2 5982 (Imaging of a Complete Sample of the Nearest Infrared Quasars) The WF/PC-2 was used to make two color observations of the nearby infrared quasar (IR QSO) MRK1014. It is believed that this IR QSO resulted from the merger of two gas-rich galaxies. The observations were executed as planned, and no anomalies were reported. 1.16 Completed GHRS/1 5950 (Spectroscopy of the Three UV Brightest Known High-Redshift Quasars) The High Resolution Spectrograph (Side One) was used to make ultraviolet observations of quasar HS-1307+46 which is know to be one of the brightest high redshift quasars. The objectives are to measure the C, N, O, and Ne abundances in the halos of early galaxies, to measure the intensity of the ionizing UV background, to measure He I, and to study the Lyman-Alpha forest. The observations were completed as scheduled, and no anomalies were reported. 1.17 Completed FOC/96 6156 (The Intergalactic HE II Absorption in PKS 1935-6792 and Q0302-003) The Faint Object Camera (f/96) was used with the prism to obtain far-ultraviolet spectra of quasar Q1935-692 to study the intergalactic absorption of He II. This is a high redshift quasar (Z=3.18) which has much less intergalactic helium absorption than other quasars at the same distance. The observations were completed as planned, and no problems were reported. 1.18 Completed WF/PC-2 6191 (WF/PC-2 Cycle 4 Calibration: UV Flats) The WF/PC-2 was used to take some ultraviolet calibrations flats using several different filters. A dark calibration image was also taken to check for possible after images. No problems were reported. 1.19 Completed FGS 6047 (Calibrating the Mass Luminosity Relation at the End of the Main Sequence) The FGSs were used to make astrometric observations of star GL748 as part of a study to determine its mass luminosity relationship. The purpose is to measure the precise parallax and relative orbit of its faint companion. The observations were completed as planned, and no problems were reported. 1.20 Completed WF/PC-2 6186 (UV Throughput) The WF/PC-2 was used to make ultraviolet photometric observations of standard star GRW+70D582 in order to measure the ultraviolet throughput of the system. This was done shortly before and after a DECON though all the UV filters in each chip. The observations were completed as planned, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 17 Successful : 17 Scheduled Reacquisitions : 23 Successful : 23 2.2 FHST Updates: Scheduled : 43 Successful : 43 2.3 Operations Notes: Throughtout the weekend, operations experienced periodic problems with the VDS communications system. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-